# Marrying Autoregressive Transformer and Diffusion with Multi-Reference Autoregression

### v1

@@ Appendix C The Euler–Maruyama Method
Appendix D More Quantitative Results
@@
Type Model #params FID ↓ IS ↑ Time ↓
Diffusion ADM [10] 554M 23.24 101.0 -
Diffusion DiT-XL/2 [33] 657M 3.04 240.8 81s
AR VQGAN [12] 227M 26.52 290.5 25s
VAR VAR-d36-s [46] 2.3B 2.63 303.2 1s
TransDiff TransDiff-L, MRAR 683M 2.51 276.6 1s
Appendix E More Qualitative Results
We also provide more visualization examples in Figure 78 and 9. As presented, Transdiff-H MRAR produces high-quality, diverse, and content-rich image samples. Furthermore, as illustrated in Figure 7, compared to VAR and MAR under similar parameter settings, Transdiff produces more realistic textures, sharper facial features, and consistent lighting. This highlights its strong potential for complex semantic tasks like human face synthesis.

### v3

@@ Appendix F The Euler–Maruyama Method
Appendix G More Quantitative Results
Type Model #params FID ↓ IS ↑ Time ↓
Diffusion ADM 554M 23.24 101.0 -
Diffusion DiT-XL/2 657M 3.04 240.8 81s
AR VQGAN 227M 26.52 290.5 25s
VAR VAR-d36-s 2.3B 2.63 303.2 1s
TransDiff TransDiff-L, MRAR 683M 2.51 276.6 1s
@@ Appendix G More Quantitative Results
Figure 7: Correlation between the number of references and FID score.
Appendix H More Qualitative Results
We also provide more visualization examples in Figure 89 and 10. As presented, Transdiff-H MRAR produces high-quality, diverse, and content-rich image samples. Furthermore, as illustrated in Figure 8, compared to VAR and MAR under similar parameter settings, Transdiff produces more realistic textures, sharper facial features, and consistent lighting. This highlights its strong potential for complex semantic tasks like human face synthesis.